# Latent Dataset Distillation with Diffusion Models

### v2

@@ Abstract
The efficacy of machine learning has traditionally relied on the availability of increasingly larger datasets. However, large datasets pose storage challenges and contain non-influential samples, which could be ignored during training without impacting the final accuracy of the model. In response to these limitations, the concept of distilling the information on a dataset into a condensed set of (synthetic) samples, namely a distilled dataset, emerged. One crucial aspect is the selected architecture (usually ConvNet) for linking the original and synthetic datasets. However, the final accuracy is lower if the employed model architecture differs from the model used during distillation. Another challenge is the generation of high-resolution images, e.g., 128x128 and higher. In this paper, we propose Latent Dataset Distillation with Diffusion Models (LD3M) that combine diffusion in latent space with dataset distillation to tackle both challenges. LD3M incorporates a novel diffusion process tailored for dataset distillation, which improves the gradient norms for learning synthetic images. By adjusting the number of diffusion steps, LD3M also offers a straightforward way of controlling the trade-off between speed and accuracy. We evaluate our approach in several ImageNet subsets and for high-resolution images (128x128 and 256x256). As a result, LD3M consistently outperforms state-of-the-art distillation techniques by up to 4.8 p.p. and 4.2 p.p. for 1 and 10 images per class, respectively.
@@ 1 Introduction
Following the idea of knowledge distillation [25], dataset distillation [63] aims to generate a small set of representative synthetic samples from the original training set. Since then, the idea of condensing large datasets into a few synthesized images, e.g., one image per class, without significantly compromising training, has rapidly gained traction. It has found applications in the privacy domain or accelerated neural architecture search [14, 60, 7, 38].
However, conventional methods for dataset distillation like Dataset Condensation [69], Distribution Matching [68], and Matching Training Trajectories [6] still face major limitations. The core problems are the generalization across architectures and high image resolutions, i.e. $128\times 128$ or $256\times 256$. For instance, models show peak performance when their architecture matches the one used for distillation. However, the performance deteriorates if other architectures are applied. These challenges arise primarily from the common practice of distilling the raw pixel values of synthetic images [8]. Optimizing pixels directly leads to datasets that overfit the network that is used during the dataset distillation [3].
To address these challenges, we propose Latent Dataset Distillation with Diffusion Models (LD3M), which integrates state-of-the-art image generation methods with dataset distillation. LD3M avoids distilling into the input space and finds better representations in latent space that can be decoded into synthetic samples. Training multiple architectures with the synthetic samples leads to higher accuracy overall, addressing the lack of generalization w.r.t. traditional distillation methods. The consistency of the latent space also leads to improved quality of high-resolution samples compared to other state-of-the-art methods.
One benefit of LD3M is that it exploits a pre-trained diffusion model out of the box. Moreover, LD3M is compatible on top of any distillation algorithm, turning LD3M into a flexible and more powerful alternative to raw distillation methods [68, 69, 6]. The general formulation of LD3M enables the use of any diffusion model, making it extensible to new developments in the field of diffusion models. Also, using the pre-trained LDM allows for a straightforward initialization of the distilled dataset. In other words, the original images are passed through the encoder to obtain their respective latent codes. This is an advantage over current latent distillation methods that use costly GAN inversions [5, 71].
We evaluate our approach using varying numbers of images per class, multiple model architectures, and varying image resolutions, all in the context of image classification. Our results show that LD3M is faster in distillation and generates datasets leading to higher accuracy on various architectures than the state-of-the-art method GLaD [58]. Overall, our paper:
@@
  • improves performance compared to the state-of-the-art method GLaD on various dataset distillation experiments.
@@ 2 Background
Parts of L3DM rely on different methods, including more primitive distillation algorithms. In this section, we briefly review the most important and discuss their relation to our method.
2.1 Dataset Distillation Algorithms for Image Classification
Distillation algorithms compress training image datasets into a small set of synthetic samples, whose size is defined by the number of Images Per Class (IPC). The synthetic images are learnable, similar to weights in neural networks. After termination of the distillation algorithm, the synthetic images constitute an expressive representation of a class in the original dataset. Most approaches initialize the synthetic images with IPC-many random images from each class as a starting point. The core idea is to use a simple ConvNet and align its training progress on the synthetic set with the progress on the original dataset [32, 36, 67]. This alignment can be expressed as an optimization process that updates the synthetic images in order to maximize training progress similarity. In this section, we describe three commonly used methods, namely Dataset Condensation [69], Distribution Matching [68], and Matching Training Trajectories [6]. Each method represents a different type of distillation algorithm: gradient matching, feature matching, and parameter matching.
Dataset Condensation (DC) ensures the alignment by deriving the gradients via a classification error [69]. It calculates the loss on real and the respective synthetic data. Next, it minimizes the distance between the two loss gradients.
@@ 2.1 Dataset Distillation Algorithms for Image Classification
Matching Training Trajectories (MTT) does not rely on the gradients obtained by calculating the classification error or the feature distance. Instead, it uses the network parameters [6]. It trains many model instantiations on the original dataset, called experts, and records the history of parameters at predetermined intervals, called expert trajectories. For dataset distillation, it samples a random set of parameters from this history at a given timestamp and trains a new network initialized with the parameters on the respective synthetic images. Finally, it minimizes the distance between the trajectory on the real dataset and the trajectory on the synthetic one. As a result, it tries to mimic the training path (trajectory of parameters) of the original dataset with the synthetic images.
@@ 2.2 Diffusion Models for Image Generation
The premise of diffusion models is to generate real-looking data by progressively adding noise to an initial sample and subsequently reversing the process by eliminating the noise. This reversal aids in approximating a complex target data distribution gradually. [43, 65, 42]. The main characteristic of diffusion models, which sets them apart from previous generative models, is their dependency on previous time steps. Through iterative refinement, a generative model keeps track of small perturbations and corrects them instead of predicting a large and challenging transformation [26, 50, 65]. These iterative steps unfold forward and backward in time. The distinct stages of the diffusion process are characterized as follows:
@@
Backward Diffusion - This stage involves the iterative denoising of the previously degraded data, effectively reversing the noise addition process and thus moving backward in time.
To reduce the computational requirements for generating high-resolution images, Rombach et al. propose to move the diffusion process towards the latent space representations of an autoencoder structure, which they called Latent Diffusion Model (LDM) [55]. Usually, a pre-trained autoencoder [18] is used to compress input data into a low-dimensional latent representation. Next, LDM uses the regular diffusion pipeline in the latent space and employs a decoder to translate the result, the diffusion-processed latent code, back to pixel space.
This application of the diffusion pipeline in latent space is essential for LD3M, as we want to optimize the latent codes that generate the synthetic images. Thus, LD3M directly exploits two main components of the LDM during dataset distillation: The denoising U-Net and the decoder to translate the latent code to image space. The encoder and embedding network for the conditioning will be used once to instantiate the initial latent and conditioning codes for each class.
@@ 2.3 Dataset Distillation with Generative Models
The Generative Latent Distillation (GLaD) incorporates a deep generative prior to the distillation process, serving as a form of regularization [8]. This is achieved by optimizing the latent codes of a pre-trained generative model instead of directly focusing on the raw pixel values. GLaD uses StyleGAN-XL [58], a large-scale GAN, and an adaptation of StyleGAN3 [29] as the pre-trained generator.
@@ 3 LD3M
In this work, we introduce a novel method called Latent Dataset Distillation with Diffusion Models (LD3M) that leverages the generative power of diffusion models for high-resolution image generation. Our work focuses on dataset distillation in the latent space of diffusion models to solve the generalization problem of unseen architectures and high-resolution images (i.e., 256x256). In more detail, we exploit a pre-trained Latent Diffusion Model (LDM) that maps the latent space to image space without fine-tuning [55, 44, 65]. Since LDM is a conditional diffusion model, the generative process contains two crucial elements: the latent codes and their associated conditioning code. Therefore, the challenge is to find the latent code and the conditioning code to generate the distilled dataset. In the following, we will explain the sampling process, which takes the latent and conditioning code as input and generates the synthetic image. Next, we describe how we initialize the latent codes before dataset distillation and, finally, how we use gradient checkpointing to save VRAM. A general overview of our approach is shown in Figure 1.
@@ 3.1 Sampling Process
We use a pre-trained LDM without fine-tuning. Thus, we focus primarily on the backward diffusion process $p$, which starts from initial state $\mathbf{z}_{T}$, usually Gaussian noise $\varepsilon\in\mathcal{N}(\mathbf{0},\mathbf{I})$, and performs the inference conditioned on $\mathbf{c}$, which is in our case the pre-trained embedding of the class label [65, 44]. The forward diffusion process is only needed to derive $\mathbf{z}_{T}$ from the learned and distilled latent representations $\mathcal{Z}$. Therefore, we can view $\mathbf{z}_{T}$ as the distorted representation of $\mathcal{Z}$, which the LDM refines. We approximate $p$ with a parameterized time-conditional process $p_{\theta}$, such that
@@
For dataset distillation, we make the embedded conditioning $\mathbf{c}$ as well as the latent code $\mathcal{Z}$ leading to the initial state $\mathbf{z}_{T}$ learnable. Since the LDM is not trained from scratch, only modifying the condition state $\mathbf{c}$ is insufficient and does not allow the LDM enough freedom to generate expressive synthetic images, which necessitates the need to make $\mathcal{Z}$ learnable [51, 34]. We will show this quantitatively in the experiments section. The initial experiments indicate that making only the conditioning $\mathbf{c}$ of the LDM learnable predominantly produces images resembling real data.
A significant challenge in making the latent code $\mathcal{Z}$, which leads to the initial state $\mathbf{z}_{T}$, learnable arises from the numerous time steps during the backward diffusion process. For instance, LDM employs $T=200$ steps during training. This extensive computation chain leads to vanishing gradients for $\mathbf{z}_{T}$, impeding the effective distillation of synthetic images [27]. We modify Equation 5 to counteract this by including residual connections in the computational graph, thereby enhancing gradient flow. Specifically, we integrate the initial state $\mathbf{z}_{T}$ into the computation of the intermediate states $\mathbf{z}_{t}$. Additionally, we systematically diminish its influence as $t$ approaches 0. This adjustment ensures an enhanced propagation of gradients in the computational graph, crucial for the generation of diverse and representative synthetic latent codes $\mathcal{Z}$:
@@
Note that this reformulation is not bound to LDMs and can be used by any diffusion model type for future work.
@@ 3.2 Initializing Latent Codes
We initialize the synthetic images with random pictures representing the respective class. More specifically, we derive the latent codes leading to the selected random images. We share this challenge with GLaD, which uses StyleGAN-XL before applying dataset distillation algorithms. GLaD uses GAN-inversion techniques to solve this problem, which is time-consuming. To obtain the latent code for a GAN, one has to solve the optimization problem $\mathcal{Z}^{*}=\arg\min_{\mathcal{Z}}\mathcal{L}\left(y,G(\mathcal{Z})\right)$. A solution can be optimization-based, learning-based, or hybrid-based, which needs careful adjustment and testing [64, 8]. Nonetheless, the solutions are time-consuming as they iteratively sample possible $\mathcal{Z}^{*}$ or require extra training.
In contrast, LD3M can initialize the latent codes straightforwardly by applying the pre-trained encoder to the random images. Furthermore, LD3M initializes the embedded condition information by applying the pre-trained embedding network. As a result, initialization with LD3M is more straightforward than GLaD and omits careful adjustments before applying dataset distillation.
@@ 3.3 Memory Saving via Checkpointing
Due to the substantial VRAM requirements for the forward pass in modern generative models, a straightforward implementation is not feasible with limited GPUs [20]. Therefore, as in GLaD [8], we employ gradient checkpointing [9].
@@
4 Experiments
We follow a similar setup as Cazenavette et al. [8]. Thus, we evaluate the cross-architecture performance for IPC=1 (MTT, DC, DM) and IP=10 (DC, DM) with image size $128\times 128$. Moreover, we run a high-resolution dataset distillation evaluation with DC and image size $256\times 256$ for IPC=1. Finally, we compare the results visually with GLaD and analyze the impact of time steps and different components of LD3M. We maintain a consistent set of hyper-parameters across experiments to guarantee a fair comparison. See supplementary materials for more details.
4.1 Setup
Evaluation Protocol. We first distill synthetic datasets using the designated algorithms and then assess the quality of the datasets by measuring the performance across unseen network architectures. To evaluate a synthetic dataset with a specific architecture, we train a new network from scratch on the distilled dataset and subsequently test its performance on the validation set. The training protocol is consistent across all networks and datasets for fair comparison. Each architecture employs a predetermined and constant initial learning rate. This entire process is replicated five times, with the report including the mean validation accuracy plus or minus one standard deviation.
Latent Diffusion Model. For the diffusion model, we use the ImageNet pre-trained LDM [55], which guarantees a fair comparison with the ImageNet pre-trained StyleGAN-XL used in GLaD [8]. We use the pre-trained autoencoder with 2×2$\times$2 × compression of the LDM without further adjustments to compress into and decompress out of the latent space. We employ for all experiments a max. time step $T=10$ for image size $128\times 128$ and $T=20$ for $256\times 256$. Besides this hyper-parameter, no adjustment to the pre-trained LDM was necessary.

### v3

@@ Abstract
Machine learning traditionally relies on increasingly larger datasets. Yet, such datasets pose major storage challenges and usually contain non-influential samples, which could be ignored during training without negatively impacting the training quality. In response, the idea of distilling a dataset into a condensed set of synthetic samples, i.e., a distilled dataset, emerged. One key aspect is the selected architecture, usually ConvNet, for linking the original and synthetic datasets. However, the final accuracy is lower if the employed model architecture differs from that used during distillation. Another challenge is the generation of high-resolution images (128x128 and higher). To address both challenges, this paper proposes Latent Dataset Distillation with Diffusion Models (LD3M) that combine diffusion in latent space with dataset distillation. Our novel diffusion process is tailored for this task and significantly improves the gradient flow for distillation. By adjusting the number of diffusion steps, LD3M also offers a convenient way of controlling the trade-off between distillation speed and dataset quality. Overall, LD3M consistently outperforms state-of-the-art methods by up to 4.8 p.p. and 4.2 p.p. for 1 and 10 images per class, respectively, and on several ImageNet subsets and high resolutions (128x128 and 256x256).
@@ 1 Introduction
Following the idea of knowledge distillation [25], dataset distillation [63] aims to generate a small set of representative synthetic samples from the original training set. Since then, the idea of condensing large datasets into a few synthesized images, e.g., one image per class, has rapidly gained traction. It has found applications in the privacy domain or accelerated neural architecture search [14, 60, 7, 37].
However, conventional methods like Dataset Condensation [69], Distribution Matching [68], and Matching Training Trajectories [6] still face major limitations. The fundamental problems are the generalization across architectures and high image resolutions, i.e., $128\times 128$ or $256\times 256$. Models show peak performance when their architecture matches the one used for distillation. Yet, the performance deteriorates if other architectures are applied. Furthermore, dataset distillation performs poorly on images with resolutions larger than $32\times 32$. This arises primarily from the common practice of distilling the raw pixel values [8]. Optimizing pixels directly leads to datasets that overfit the network used during the distillation process [3].
To address these challenges, we propose Latent Dataset Distillation with Diffusion Models (LD3M), which integrates state-of-the-art image generative models, namely diffusion models, with dataset distillation. LD3M avoids distilling into the pixel space and finds better representations in latent space that can be decoded into synthetic images. Training multiple architectures with the synthetic samples leads to significantly higher accuracy overall, addressing the lack of generalization of traditional distillation methods. The consistency of the latent space also allows for the improved quality of higher-resolution samples compared to other state-of-the-art methods.
One benefit of LD3M is that it exploits a pre-trained diffusion model out of the box. Moreover, LD3M is compatible with any distillation algorithm, turning LD3M into a flexible and more powerful alternative to raw distillation methods [68, 69, 6]. The general formulation of LD3M enables the use of any diffusion model, making it easily extensible and adaptable to new developments in the field of diffusion models. Also, using a pre-trained LDM allows for a straightforward initialization of the distilled dataset. In other words, the original images are passed through the encoder to obtain their respective latent codes. This is an advantage over current GAN-based dataset distillation methods like GLaD [8] that use costly GAN inversions [5, 72].
We evaluate our approach on varying numbers of images per class, multiple model architectures, and different image resolutions, all in the context of image classification. Our results show that LD3M distills notably faster and generates datasets of much higher quality for training various architectures than the state-of-the-art latent dataset distillation method GLaD [57]. Overall, our work:
@@
  • significantly improves the quality of synthetic datasets compared to the state-of-the-art method GLaD in various dataset distillation experiments.
@@ 2 Background
This section briefly reviews the most important distillation algorithms and discusses their relation to LD3M.
2.1 Dataset Distillation Algorithms
Distillation algorithms compress datasets into a small set of synthetic samples, the size of which is defined by the number of Images Per Class (IPC). Initially, the synthetic images are learnable, similar to weights in neural networks. After the distillation, the synthetic images constitute an expressive representation of a class in the original dataset. Most approaches initialize the synthetic images with IPC-many random images from each class. The core idea is then to use a simple ConvNet and to align its training progress on the synthetic set with the progress on the original dataset [31, 35, 67]. This alignment can be expressed as an optimization process that updates the synthetic images to maximize similarity in training progress. In this section, we describe three commonly used methods, namely Dataset Condensation [69], Distribution Matching [68], and Matching Training Trajectories [6]. Each method represents a different type of distillation algorithm: gradient matching, feature matching, and parameter matching.
Dataset Condensation (DC) ensures alignment by deriving the gradients via a classification error [69]. It calculates the loss on real and the respective synthetic data. Next, it minimizes the distance between the two loss gradients.
@@ 2.1 Dataset Distillation Algorithms
Matching Training Trajectories (MTT) does not rely on the gradients obtained by calculating the classification error or the feature distance. Instead, it uses the network, i.e., ConvNet, parameters [6]. It trains multiple model instantiations on the original dataset, called experts, and records the history of parameters at predetermined intervals, called expert trajectories. For dataset distillation, it samples a random set of parameters from this history at a given timestamp. Next, it trains a new network initialized with the parameters on the respective synthetic images. Finally, it minimizes the distance between the trajectory on the real dataset and the trajectory on the synthetic one. As a result, it tries to mimic the original dataset’s training path (trajectory of parameters) with the synthetic images.
@@ 2.2 Diffusion Models for Image Generation
The premise of diffusion models is to generate real-looking data by progressively adding noise to an initial sample and subsequently reversing the process by eliminating the noise. This reversal aids in approximating a complex target data distribution gradually [42, 65, 41, 43]. The main characteristic of diffusion models, which sets them apart from previous generative models, is their dependency on previous time steps. Through iterative refinement, the generative model keeps track of small perturbations and corrects them instead of predicting a large and challenging transformation [26, 49, 65]. These iterative steps unfold forward and backward in time:
@@
Backward Diffusion - Iteratively denoise the previously degraded data, effectively reversing the noise addition process and thus moving backward in time.
To reduce the computational requirements for generating high-resolution images, Rombach et al. propose to move the diffusion process towards the latent space representations of an autoencoder structure, which they called Latent Diffusion Model (LDM) [54]. Usually, a pre-trained autoencoder [18] is used to compress input data into a low-dimensional latent code. Next, LDM uses the diffusion pipeline in the latent space and employs a decoder to translate the result, the processed latent code, back to pixel space.
This application of diffusion in latent space is essential for LD3M, as we want to optimize the latent codes that generate synthetic images. LD3M directly exploits two main components of the LDM: the denoising U-Net and the decoder, which translates the latent code to image space. The encoder will be used once to instantiate each class’s initial latent codes.
@@ 2.3 Dataset Distillation with Generative Models
The Generative Latent Distillation (GLaD) incorporates a deep generative prior to the distillation process, serving as a form of regularization [8]. This is achieved by optimizing the latent codes of a pre-trained generative model instead of directly focusing on the raw pixel values. GLaD uses StyleGAN-XL [57], a large-scale and resource-costly GAN as the pre-trained generator.
Like most distillation methods, it initializes its synthetic samples with latent codes of real images. Distilling the latent codes of an image in StyleGAN-XL is similar to a pseudo-inversion task, connecting to the broader research domain of GAN inversion [5, 72]. The central concept involves mapping an actual input image into GAN latent codes, where selecting a particular latent code balances expressiveness and reconstruction fidelity, as evidenced in prior research [62, 74]. The cost-intensive GAN inversion process is a significant drawback compared to our method LD3M, as LD3M can initialize the initial latent codes by straightforwardly applying a pre-trained encoder.
@@ 3 LD3M
In this work, we introduce a novel method called Latent Dataset Distillation with Diffusion Models (LD3M) that leverages the generative power of diffusion models for high-resolution image distillation. As in GLaD, we focus on dataset distillation in the latent space to solve the generalization problem of unseen architectures and high-resolution images, i.e., 256x256. Nonetheless, instead of using a large StyleGAN-XL as a backbone, we will refer to a lighter and more expressive generative model, namely a diffusion model [59, 44, 13]. In more detail, we exploit a pre-trained Latent Diffusion Model (LDM) that maps the latent space to image space without fine-tuning [54, 43, 65]. Since LDM is a conditional diffusion model, the generative process contains two crucial elements: the latent codes and their associated conditioning code. Therefore, the challenge is to find the latent code and the conditioning code to generate the distilled dataset. In the following, we will explain the sampling process, which takes the latent and conditioning code as input and generates the synthetic image. Next, we describe how we initialize the latent codes before dataset distillation and, finally, how we use gradient checkpointing to reduce VRAM consumption. A general overview of our approach is shown in Figure 1.
@@ 3.1 Sampling Process
We use a pre-trained LDM without fine-tuning. Thus, we focus primarily on the backward diffusion process $p$, which starts from an initial state $\mathbf{z}_{T}$, usually Gaussian noise $\varepsilon\in\mathcal{N}(\mathbf{0},\mathbf{I})$, and performs the inference conditioned on $\mathbf{c}$, which is in our case a pre-trained embedding of the class label [65, 43]. The forward diffusion process is only needed to derive $\mathbf{z}_{T}$ from the learned and distilled latent representations $\mathcal{Z}$. Therefore, we can view $\mathbf{z}_{T}$ as the distorted representation of $\mathcal{Z}$, which the LDM refines. We approximate $p$ with a parameterized time-conditional process $p_{\theta}$, such that
@@
For dataset distillation, we make the embedded conditioning $\mathbf{c}$ as well as the latent code $\mathcal{Z}$ leading to the initial state $\mathbf{z}_{T}$ learnable. Since the LDM is not trained from scratch, only modifying the condition state $\mathbf{c}$ is insufficient and does not allow the LDM enough freedom to generate expressive synthetic images, which necessitates the need to make $\mathcal{Z}$ learnable [50, 33]. We will show this quantitatively in the experiments section. Our initial experiments indicate that making only the conditioning $\mathbf{c}$ of the LDM learnable predominantly produces images resembling real data.
A significant challenge in making the latent code $\mathcal{Z}$, which leads to the initial state $\mathbf{z}_{T}$, learnable arises from the numerous time steps during the backward diffusion process. For instance, LDM employs $T=200$ steps during training. This extensive computation chain leads to vanishing gradients for $\mathbf{z}_{T}$, impeding an effective distillation [27]. We modify Equation 5 to counteract this by including residual connections in the computational graph, thereby enhancing gradient flow. Specifically, we integrate the initial state $\mathbf{z}_{T}$ into the computation of the intermediate states $\mathbf{z}_{t}$. Additionally, we systematically diminish its influence as $t$ approaches 0. This adjustment ensures an enhanced gradient propagation crucial for the generation of diverse and representative synthetic latent codes $\mathcal{Z}$:
@@
Note that this reformulation is not bound to LDMs and can be used by any diffusion model type for future work. However, we focused on LDM because of several advantages: (a) there exists a pre-trained model, (b) it is foundational for diffusion models (especially in exploiting latent space), and (c) it serves as a proof of concept.
@@ 3.2 Initializing Latent Codes
We initialize the synthetic images with random samples representing the respective class. More specifically, we derive the latent codes leading to the selected random images. We share this challenge with GLaD, which uses StyleGAN-XL before applying dataset distillation algorithms. GLaD uses GAN-inversion techniques to solve this problem, which is time-consuming. To obtain the latent code for a GAN, one has to solve the optimization problem $\mathcal{Z}^{*}=\arg\min_{\mathcal{Z}}\mathcal{L}\left(y,G(\mathcal{Z})\right)$. A solution can be optimization-based, learning-based, or hybrid-based, all needing careful adjustments and testing [64, 8]. Moreover, finding solutions is time-consuming as an iterative sampling of possible $\mathcal{Z}^{*}$ or extra training is required.
In contrast, LD3M can initialize the latent codes straightforwardly by applying the pre-trained encoder to the random images. Furthermore, LD3M initializes the embedded condition information by applying the pre-trained embedding network. As a result, initialization with LD3M is more straightforward and computationally more efficient than GLaD and omits careful adjustments before applying dataset distillation.
@@ 3.3 Memory Saving via Checkpointing
Due to the substantial VRAM requirements for the forward pass in modern generative models, a straightforward use is not feasible with limited GPUs [20]. Thus, like GLaD [8], we employ gradient checkpointing [9].
@@
4 Experiments
We follow a similar setup as Cazenavette et al. [8] and evaluate the cross-architecture performance for IPC=1 (MTT, DC, DM) and IP=10 (DC, DM) with image size $128\times 128$. Furthermore, we run a high-resolution dataset distillation evaluation with DC and image size $256\times 256$ for IPC=1 as well as a low-resolution dataset distillation evaluation on CIFAR-10. Finally, we compare the results visually with GLaD and analyze the impact of different latent and conditioning code initializations, various time steps and different components of LD3M. We maintain a consistent set of hyper-parameters across experiments to guarantee a fair comparison (more details in the appendix).
4.1 Setup
Evaluation Protocol. We first distill synthetic datasets using the designated algorithms and then assess the quality of the datasets by measuring the performance across unseen network architectures. To evaluate a synthetic dataset with a specific architecture, we train a new network from scratch on the distilled dataset and subsequently test its performance on the validation set. This entire process is replicated five times, with the report including the mean validation accuracy plus or minus one standard deviation.
Latent Diffusion Model. For the diffusion model, we use the ImageNet pre-trained LDM [54], which guarantees a fair comparison with the ImageNet pre-trained StyleGAN-XL used in GLaD [8]. We use the pre-trained autoencoder with 2×2$\times$2 × compression of the LDM without further adjustments to compress into and decompress out of the latent space. We employ for all experiments a max. time step $T=10$ for image size $128\times 128$ and $T=20$ for $256\times 256$. Besides this hyper-parameter, no adjustment to the pre-trained LDM was necessary. In addition, we want to highlight that using a pre-existing diffusion model is not a limitation but an advantage, showing that a pre-trained diffusion model can be used out of the box (and empirically more consistent as GLaD w.r.t. various experiments).
4.2 Results